OpenStudy (anonymous):

Is the sequence 3, 12, 36, ... a geometric sequence?

OpenStudy (ash2326):

A geometric sequence is of the form \[a, ar, ar^2, ar^3...\] Two consecutive terms are in the same ratio Let's check if the sequence satisfies this \[\frac{12}{3}=\frac{36}{12}\] \[4 \ne 3\] So this ain't a geometric series

OpenStudy (anonymous):

In a geometric sequence you get the (n+1)'th term by multiplying the n'th term by the common ratio r.To test if 3,12,36 ... are the terms of a geometric sequence we calculate the common ratio for 3 and 12 ,and 12 and 36. Thus: r = 12/3 = 4 and r = 36/12 = 3 Since 4 does not equal 3 there is not a constant common ratio and hence this is not a geometric sequence

OpenStudy (anonymous):

oh okay thanks guys!! :)

OpenStudy (anonymous):

yw